A man rides a bus part of the way to work and walk the rest. the bus speed is 36mph and he walks 4mph the distance from home is 15 mi total time is 2 hr.
how far does he walk and how far does he ride

Answer by Alan3354(69443)

A man rides a bus part of the way to work and walk the rest. the bus speed is 36mph and he walks 4mph the distance from home is 15 mi total time is 2 hr.
how far does he walk and how far does he ride
---------------
d = r*t
t = d/r
----
b = bus distance
w = walking distance
---
b/36 + w/4 = 2 Eqn 1
b + w = 15 Eqn 2
------
b + 9w = 72 (Eqn 1*36)
b + w = 15
----------------------------------- subtract
8w = 57
w = 57/8
etc